OpenStudy (anonymous):

Solve the following equation using the Bernoulli substitution method: y' = εy - δy³, ε>0 & δ>0

OpenStudy (anonymous):

I got it down to 2v' + εv = δ but I'm not sure if that's correct or not.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

it's correct

OpenStudy (anonymous):

So I found the integrating factor to be E^(εx) which leads the equation to E^(εx)∗v=∫(E^(εx))∗δdx delta can't be a constant right? So I would need to do a partial integration but this doesn't seem right...

OpenStudy (anonymous):

\[e ^{\int\limits_{}^{}\epsilon } ?\]

OpenStudy (anonymous):

aren't delta and epsilon supposed to be functions of x?

OpenStudy (anonymous):

or just constants?

OpenStudy (anonymous):

I'm not sure at this point, the constraints for the problem is delta >0 and epsilon > 0 :\

OpenStudy (anonymous):

constants I guess... weird.. if that's so then you're right up to...

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

should have: \[v = e^{-\epsilon x} * \int\limits_{ }^{ } e^{\epsilon x} * \delta * dx\]

OpenStudy (anonymous):

yeah you got that

OpenStudy (anonymous):

simplifies nicely

OpenStudy (anonymous):

\[\delta / \epsilon ?\]

OpenStudy (anonymous):

for v? is that what you got?

OpenStudy (anonymous):

I got... (δ/ϵ) + (C/e^(δx))

OpenStudy (anonymous):

right good call, makes no sense otherwise (y would be a constant lol)

OpenStudy (anonymous):

forgot my C

OpenStudy (anonymous):

Ah alright, after some simplifying I got it to match the answer

OpenStudy (anonymous):

\[y = v ^{-1/2}\] so you're done

OpenStudy (anonymous):

Thanks for your guidance :)

OpenStudy (anonymous):

Thank you for yours, haven't played with these for years; didn't think anyone taught them anymore..

OpenStudy (anonymous):

Heh, yeah I'm taking O.D.E. right now and the professor is going at a breakneck pace haha

OpenStudy (anonymous):

good times... :) gl2u!

OpenStudy (anonymous):

Thanks :p